In the study of urinary tract infection, the sample and proportion in the group of

fosfomycin / trometamol was 25, 0.92 and trimethoprim / sulfamethoxazole was

16, 0.61, respectively.

a) Find the difference in proportions. (0.5)

4

b) Standard error of difference in proportions. (0.5)

c) Compute a 95% confidence interval of difference in proportions


1
Expert's answer
2022-01-31T17:41:15-0500

Given,

population 1

"n_1=25,\\\hat p_1 =0.92"

population 2

"n_2=

16,\\\hat p_2= 0.61,"


"a)"

The difference in proportion is "\hat p_1-\hat p_2=0.92-0.61=0.31"


"b)"

The standard error of difference in proportions is given as,

"SE(\hat p_1-\hat p_2)=\sqrt{{\hat p_1\times (1-\hat p_1)\over n_1}+{\hat p_2\times (1-\hat p_2)\over n_2}}=\sqrt{{0.92\times0.08\over 25}+{0.61\times0.39\over16}}=\sqrt{0.002944+0.01486875}=0.13346441"

Therefore, the standard error of difference in proportions is 0.13346441


"c)"

A 95% confidence interval of difference in proportions is given as,

"CI=(\hat p_1-\hat p_2)\pm Z_{\alpha\over2}SE(\hat p_1-\hat p_2)" where "Z_{\alpha\over2}" is the standard normal table value at "\alpha=0.05". Thus "Z_{0.05\over2}=Z_{0.025}=1.96"

Hence,

"CI=0.31\pm(1.96\times 0.13346441)=0.31\pm0.26159024"

Therefore, a 95% confidence interval for the difference in proportions is, "(0.0484,0.5716)"